OpenStudy (anonymous):

Graph an exponential function f(x) whose domain is all real numbers and is such that: f(x) is decreasing over it's entire domain, f(x)=3 as x approaches - infinity, f(x)=-infinity as x approaches +infinity. Its a little confusing on these limits, thanks.

OpenStudy (anonymous):

The range of the exponent function, \[f(x)=e^x\] is \[(0,\infty)\] . The range of \[f(x)=-e^{x}\] is \[(0,-\infty)\] if you add a constant, k, the function becomes \[f(x)=-e^x+k\] with a range of \[(k,-\infty)\] k=3, \[f(x)=-e^x+3\]

OpenStudy (tkhunny):

If you build it, you will see it. f(x) = e^x Domain is good, but -infty ==> 0 and +infty ==> +infty We're upside down. f(x) = -e^x Domain is good and +infty ==> -infty, but -infty ==> 0 We're three low. f(x) = 3-e^x Domain is good, -infty ==> 3 and +infty ==> -infty Done.

OpenStudy (anonymous):